# Joint Problems in Learning Multiple Dynamical Systems

### v2

@@ Abstract
Clustering of time series is a well-studied problem [58, 32], with applications ranging from quantitative, personalized models of metabolism obtained from metabolite concentrations to state discrimination in quantum information theory. We consider a variant, where given a set of trajectories and a number of parts, we jointly partition the set of trajectories and learn linear dynamical system (LDS) models for each part, so as to minimize the maximum error across all the models. We present globally convergent methods and EM heuristics, accompanied by promising computational results.
@@ 1 Introduction
The task of grouping similar time series based on their dynamic patterns has attracted significant attention due to its applications ranging from studying mobility patterns [44] to improving Apple Maps [13], through quantitative, personalized models of metabolism obtained from metabolite concentrations, all the way to state discrimination problems in quantum information theory [41].
@@
Considering the above contributions with previous work produced by [37, 28, 43], a pioneering method can be created that holds following properties:
  • Mimic temporal dynamics in the data and produces the related, interpretable features, unlike methods relying on deep learning.
  • Handle multiple time lags within the time series without losing its periodicity, unlike methods based on Fourier coefficients [30].
  • Not necessary to make any linear-algebraic assumptions, unlike subspace methods [2, 19] that assume the existence of a subspace, which might not necessarily be true. The assumption that the time series were generated by an LDS does not cause any loss of generality.
  • Avoid of making any assumptions on the dimension of the hidden state, unlike previous methods utilizing LDS [37, 38, 5, 43].
  • Improves upon previously proposed methods based on Fast Fourier Transform (FFT) [30] and Dynamic Time Warping (DTW) [53, 23, 54] in practical experiments of Section 5.
@@ 2.2 Clustering with LDS Assumptions
The problem of clustering of time series is relevant in many fields, including [1, 62] applications in Bioinformatics, Multimedia, Robotics, Climate, and Finance. There are a variety of existing methods, including those based on (Fast) Discrete Fourier Transforms (FFTs), Wavelet Transforms, Discrete Cosine Transformations, Singular Value Decomposition, Levenshtein distance, and Dynamic Time Warping (DTW). We compare our method with the FFT- and DTW-based methods below.
@@
Also at ICML 2023, [31] presented learning of mixtures of Hidden Markov Models and Markov Chains. The method estimates subspaces first, then uses spectral clustering and MDP transition matrices estimation step followed by the possibility to classify new trajectories. The paper provides performance guarantees for each step.
Still at the same conference, [71] consider the learning of a single LDS across all trajectories and fine-tuning the target policy for some subsets thereof. The main contribution of [71] is a bound on the imitation gap over trajectories generated by the learned target policy.
Finally, there is a notable arxiv pre-print. A joint problem related to [71] is being considered by [43], who propose to jointly estimate transition matrices of multiple linear time-invariant dynamic systems, where the transition matrices are all linear functions of some shared unknown base matrix. Also, as opposed to our method, the state of the system is assumed to be fully observable.
@@ 2.3 Mixed-Integer Programming (MIP)
While MINLP problems may seem too general a model for our joint problem, notice that the NP-hardness and inapproximability results discussed in Section 2.1 suggest that this may be the appropriate framework. For bounded variables, standard branch-and-bound-and-cut algorithms run in finite time. Both in theory – albeit under restrictive assumptions, such as in [18] – and in practice, the expected runtime is often polynomial. In the formulation of the next section, $f,g_{i}$ are trilinear, and various monomial envelopes have been considered and implemented in global optimization solvers such as BARON [52], SCIP [7], and Gurobi. In our approach, we consider a mixed-integer programming formulation of a piecewise polyhedral relaxation of a multilinear term using its convex-hull representation.
@@ 2.4 Non-Commutative Polynomial Optimization (NCPOP)
To extend the search for global optima from a fixed finite-dimensional state to an operator in an unknown dimension, we formulate the problem as a non-commutative polynomial optimization problem (NCPOP), cf. [11]. In contrast to traditional scalar-valued, vector-valued, or matrix-valued optimization techniques, the variables considered in NCPOP are operators, whose dimensions are unknown a priori.
@@
Under the Archimedean assumption, such that the tuple of operators $X$ are bounded, one can utilize the Sums of Squares theorem of [25] and [42] to derive semidefinite programming (SDP) relaxations of the Navascules-Pironio-Acin (NPA) hierarchy [46, 49]. There are also variants [61, 60, 59] that exploit various forms of sparsity.
@@ 3.1 Least-Squares Objective Function
$\min_{\begin{subarray}{c}f^{0}_{t},\mathbf{F}_{0},\mathbf{G}_{0},\upsilon^{0}_{t},\omega^{0}_{t},\phi_{0}^{0},\\ f^{1}_{t},\mathbf{F}_{1},\mathbf{G}_{1},\upsilon^{1}_{t},\omega^{1}_{t},\phi_{0}^{1}\\ l_{i}\end{subarray}}\sum_{i=1}^{N}\sum_{t=1}^{T}\|X^{i}_{t}-f^{l_{i}}_{t}\|_{2}^{2}\\ +\sum_{c\in\{0,1\}}\sum_{t=1}^{T}\left[\|\upsilon_{t}^{c}\|_{2}^{2}+\|\omega_{t}^{c}\|_{2}^{2}\right]\\ \omega_{t}^{c}\sim N(0,\mathbf{\Sigma}_{H}^{c}),\upsilon_{t}^{c}\sim N(0,\mathbf{\Sigma}_{O}^{c}),$ (8)
@@ 3.2 Feasible Set in State Space
We can also add redundant constraints for all $i=1,2,\ldots,N$:
$l_{i}\phi^{1}_{t}$ $=l_{i}(\mathbf{G}_{1}\phi^{1}_{t-1}+\omega^{1}_{t}),$ $\forall t=2,\dots,T,$
$l_{i}f^{1}_{t}$ $=l_{i}(\mathbf{F}^{{}^{\prime}}_{1}\phi^{1}_{t}+\upsilon^{1}_{t}),$ $\forall t=1,\dots,T,$
$(1-l_{i})\phi^{0}_{t}$ $=(1-l_{i})(\mathbf{G}_{0}\phi^{0}_{t-1}+\omega^{0}_{t}),$ $\forall t=2,\dots,T,$
$(1-l_{i})f^{0}_{t}$ $=(1-l_{i})(\mathbf{F}^{{}^{\prime}}_{0}\phi^{l}_{t}+\upsilon^{0}_{t}),$ $\forall t=1,\dots,T,$
@@ 3.3 Variants and Guarantees
  • bound the magnitude of the process noise $\omega^{c}$ and observation noise $\upsilon^{c}$, or other shape constraints thereupon – or not.
@@ 5.1 Methods and Solvers
Our formulation for MINLP with an indicator function (MIP-IF) uses equations (9) subject to (1012). We specify the dimension of system matrix $\mathbf{G}$, i.e., $n$ as the hyperparameter. For every data set, we perform $50$ experiments, in which different random seeds are employed to initialize the indicator function in each iteration. These MIP instances are solved via Bonmin [^1] https://www.coin-or.org/Bonmin/ [9].
@@
Iterated EM Heuristic clustering is presented in Algorithm 1. As in MIP-IF, the dimension, $n$, of the system matrix is required. In each iteration, the cluster partition is randomly initialized and we conduct $50$ trials with various random seeds for every dataset.
The discussion of runtime As a subroutine of the EM Heuristic, the LDS identification in equation (35) is solved via Bonmin, with runtime presented in the center subplot of Figure 3. When the dimensions of system matrices are not assumed, the LDS identification becomes an NCPOP, and the runtime increases exponentially if NPA hierarchy [46, 49] is used to find the global optimal solutions, but stays relatively modest if sparity exploit variants [61, 60, 59] are used.
@@
  • Dynamic Time Warping (DTW) is used for measuring similarity between given time series [53, 23]. We utilize K-means on DTW distance with tslearn [55] package.
@@
Figure 2: $F_{1}$ scores of proposed methods compared to baselines. The results are based on data sampled from ECG5000 with various time window $T$ chosen from $\{30,60,90,120,140\}$. For MIP-IF and EM Heuristic, $F_{1}$ scores are generated with hidden state dimension $n=3$. A higher $F_{1}$ score indicates better clustering performance.
Figure 3: left and center: $F_{1}$ scores and runtimes of MIP-IF and EM Heuristic with various hidden state dimensions $n$ chosen from $\{2,3,4\}$ and time window $T$ chosen from $\{30,60,90,120,140\}$ respectively. right: Runtimes of EM Heuristic using NCPOP compared with the method requiring specific dimension $n$ of hidden state. $n$ is chosen from $\{2,3,4\}$.
@@ 5.3 Experiments on Real-world Data
The electrocardiogram (ECG) data ECG5000 [16] is a common dataset for evaluating methods for ECG data, which has also been utilized by other papers [28] on clustering with LDS assumptions. The original data comes from Physionet [4, 22] and contains a $20$-hour-long ECG for congestive heart failure. After processing, ECG5000 includes 500 sequences, where there are $292$ normal samples and $208$ samples of four types of heart failure. Each sequence contains a whole period of heartbeat with $140$ time stamps.
@@
We randomly sample two clusters from normal sequences and one type of abnormal sequences respectively. As the entire period of time series data is not always available, we also extract subsequences with various lengths of time window $T$ chosen from $\{30,60,90,120,140\}$ to test the clustering performance.
In Figure 2, with the assumption of the hidden state dimension $n=3$, we implement all methods for $50$ runs at each length of time window. Our methods exhibit competitive performance relative to FFT and DTW when $T=140$. When the time window decreases, the performance of the baselines significantly deteriorates, while our methods maintain a higher level of robustness.
@@
Finally, when the dimension $n$ of the hidden state is not assumed, the subproblem of the EM heuristic becomes an NCPOP. For the implementation with such an assumption, we construct NCPOP using ncpol2sdpa 1.12.2222https://ncpol2sdpa.readthedocs.io/en/stable/ [65]. Subsequently, the relaxation problem is solved by Mosek 10.1333https://www.mosek.com/ [45]. Noted that the execution time of NCPOP escalates rapidly as the trajectory length $T$ grows, we test NCPOP and compare its performance with the aforementioned EM Heuristic method with $T\in\{10,20,30\}$. For comparison, we use pyomo [^4] https://www.pyomo.org/ [12] to construct the model and solve the problem with Bonmin [^5] https://www.coin-or.org/Bonmin/ [9], as above, with the dimension $n$ from $\{2,3,4\}$. The overall performance is illustrated in Table 1. NCPOP demonstrates the best performance in terms of the $F_{1}$ score when $T=10$ and $30$. However, as shown in the right subplots of Figure 3, the runtime of NCPOP grows significantly as the length $T$ of the trajectory increases.

### v3

@@ Abstract
Clustering of time series is a well-studied problem [52, 29], with applications ranging from quantitative, personalized models of metabolism obtained from metabolite concentrations to state discrimination in quantum information theory. We consider a variant, where given a set of trajectories and a number of parts, we jointly partition the set of trajectories and learn linear dynamical system (LDS) models for each part, so as to minimize the maximum error across all the models. We present globally convergent methods and EM heuristics, accompanied by promising computational results. The key highlight of this method is that it does not require a predefined hidden state dimension but instead provides an upper bound. Additionally, it offers guidance for determining regularization in the system identification.
@@ 1 Introduction
The task of clustering similar time series based on their dynamic patterns has attracted significant attention due to its applications ranging from studying mobility patterns [40] to improving Apple Maps [12], through quantitative, personalized models of metabolism obtained from metabolite concentrations, all the way to state discrimination problems in quantum information theory [38].
@@ 2.2 Clustering with LDS Assumptions
The problem of clustering of time series is relevant in many fields, including [1, 56] applications in Bioinformatics, Multimedia, Robotics, Climate, and Finance. There are a variety of existing methods, including those based on (Fast) Discrete Fourier Transforms (FFTs), Wavelet Transforms, Discrete Cosine Transformations, Singular Value Decomposition, Levenshtein distance, and Dynamic Time Warping (DTW). We compare our method with the FFT- and DTW-based methods.
@@ 2.3 Mixed-Integer Programming (MIP)
While MINLP problems may seem too general a model for our joint problem, notice that the NP-hardness and inapproximability results discussed in Section 2.1 suggest that this may be the appropriate framework. For bounded variables, standard branch-and-bound-and-cut algorithms run in finite time. Both in theory – albeit under restrictive assumptions, such as in [17] – and in practice, the expected runtime is often polynomial. In the formulation of the next section, $f,g_{i}$ are trilinear, and various monomial envelopes have been considered and implemented in global optimization solvers such as BARON [48], SCIP [6], and Gurobi  [^1] https://www.gurobi.com/ [22]. In our approach, we consider a mixed-integer programming formulation of a piecewise polyhedral relaxation of a multilinear term using its convex-hull representation.
@@ 2.4 Non-Commutative Polynomial Optimization (NCPOP)
To extend the search for global optima from a fixed finite-dimensional state to an operator in an unknown dimension, we formulate the problem as a non-commutative polynomial optimization problem (NCPOP), cf. [46, 10]. In contrast to traditional scalar-valued, vector-valued, or matrix-valued optimization techniques, the variables considered in NCPOP are operators, whose dimensions are unknown a priori.
@@
Under the Archimedean assumption, such that the tuple of operators $X$ are bounded, one can utilize the Sums of Squares theorem of [24] and [39] to derive semidefinite programming (SDP) relaxations of the Navascules-Pironio-Acin (NPA) hierarchy [45]. There are also variants [55, 54, 53] that exploit various forms of sparsity.
@@ 3.1 Least-Squares Objective Function
$\min_{\begin{subarray}{c}l_{i}\in\{0,1\}\end{subarray}}\sum_{i=1}^{N}\sum_{t=1}^{T}\|X^{i}_{t}-f^{l_{i}}_{t}\|_{2}^{2}+\sum_{c\in\{0,1\}}\sum_{t=1}^{T}\left[\|\upsilon_{t}^{c}\|_{2}^{2}+\|\omega_{t}^{c}\|_{2}^{2}\right]\\ \omega_{t}^{c}\sim N(0,\mathbf{\Sigma}_{H}^{c}),\upsilon_{t}^{c}\sim N(0,\mathbf{\Sigma}_{O}^{c}),$ (8)
@@ 3.3 Variants and Guarantees
  • bound the magnitude of the process noise $\omega^{c}$ and observation noise $\upsilon^{c}$, or other shape constraints thereupon.
@@ 5.1 Methods and Solvers
Our formulation for MINLP with an indicator function (MIP-IF) uses equations (9) subject to (1012). We specify the dimension of system matrix $\mathbf{G}$, i.e., $n$ as the hyperparameter. For every data set, we perform $50$ experiments, in which different random seeds are employed to initialize the indicator function in each iteration. These MIP instances are solved via Bonmin [^2] https://www.coin-or.org/Bonmin/ [8] and Gurobi  [^3] https://www.gurobi.com/ [22].
@@
Iterated EM Heuristic clustering is presented in Algorithm 1. As in MIP-IF, the upper bound of dimension, $n$, of the system matrix is required. In each iteration, the cluster partition is randomly initialized and we conduct $50$ trials with various random seeds for every dataset.
The discussion of runtime As a subroutine of the EM Heuristic, the LDS identification in equation (35) is solved via Bonmin, with runtime presented in the center subplot of Figure 3. When the dimensions of system matrices are not assumed, the LDS identification becomes an NCPOP, and the runtime increases exponentially if NPA hierarchy [45] is used to find the global optimal solutions, but stays relatively modest if sparity exploit variants [55, 54, 53] are used.
@@
  • Dynamic Time Warping (DTW) is used for measuring similarity between given time series [49, 21]. We utilize K-means on DTW distance with tslearn [JMLR:v21:20-091] package.
@@
Figure 2: $F_{1}$ scores of proposed methods compared to baselines. The results are based on data sampled from ECG5000 with various time window $T$ chosen from $\{30,60,90,120,140\}$. A higher $F_{1}$ score indicates better clustering performance.
Figure 3: left and center: $F_{1}$ scores and runtimes of MIP-IF and EM Heuristic with various time window $T$ chosen from $\{30,60,90,120,140\}$ respectively. The $F_{1}$ score improves as the time window increases. right: Runtimes of EM Heuristic using NCPOP compared with the method requiring specific dimension $n$ of hidden state. $n$ is chosen from $\{2,3,4\}$.
@@ 5.3 Experiments on Real-world Data
The test on electrocardiogram (ECG) data gives an inspiring application on guiding cardiologist’s diagnosis and treatment [43]. The ECG data ECG5000 [15] is a common dataset for evaluating methods for ECG data, which has also been utilized by other papers [27] on clustering with LDS assumptions. The original data comes from Physionet [3, 20] and contains a $20$-hour-long ECG for congestive heart failure. After processing, ECG5000 includes 500 sequences, where there are $292$ normal samples and $208$ samples of four types of heart failure. Each sequence contains a whole period of heartbeat with $140$ time stamps.
@@
We randomly sample two clusters from normal sequences and one type of abnormal sequences respectively. As the entire period of time series data is not always available, we also extract subsequences with various lengths of time window $T$ chosen from $\{30,60,90,120,140\}$ to test the clustering performance. In Figure 2, with the assumption of the upper bound of hidden state dimension is $n=5$, we implement all methods for $50$ runs at each length of time window. Our methods exhibit competitive performance relative to FFT and DTW when $T=140$. When the time window decreases, the performance of the baselines significantly deteriorates, while our methods maintain a higher level of robustness.
@@
Finally, when the dimension $n$ of the hidden state is not assumed, the subproblem of the EM heuristic becomes an NCPOP. For the implementation with such an assumption, we construct NCPOP using ncpol2sdpa 1.12.2444https://ncpol2sdpa.readthedocs.io/en/stable/ [59]. Subsequently, the relaxation problem is solved by Mosek 10.1555https://www.mosek.com/ [41]. Noted that the execution time of NCPOP escalates rapidly as the trajectory length $T$ grows, we test NCPOP and compare its performance with the aforementioned EM Heuristic method with $T\in\{10,20,30\}$. For comparison, we use pyomo [^6] https://www.pyomo.org/ [11] to construct the model and solve the problem with Bonmin [^7] https://www.coin-or.org/Bonmin/ [8], as above, with the dimension $n$ from $\{2,3,4\}$. The overall performance is illustrated in Table 1. NCPOP demonstrates the best performance in terms of the $F_{1}$ score when $T=10$ and $30$. However, as shown in the right subplots of Figure 3, the runtime of NCPOP grows significantly as the length $T$ of the trajectory increases.